# Estimating network-mediated causal effects via principal components network regression

### v3

@@ Abstract
We develop a method to decompose causal effects on a social network into an indirect effect mediated by the network, and a direct effect independent of the social network. To handle the complexity of network structures, we assume that latent social groups act as causal mediators. We develop principal components network regression models to differentiate the social effect from the non-social effect. Fitting the regression models is as simple as principal components analysis followed by ordinary least squares estimation. We prove asymptotic theory for regression coefficients from this procedure and show that it is widely applicable, allowing for a variety of distributions on the regression errors and network edges. We carefully characterize the counterfactual assumptions necessary to use the regression models for causal inference, and show that current approaches to causal network regression may result in over-control bias. The structure of our method is very general, so that it is applicable to many types of structured data beyond social networks, such as text, areal data, psychometrics, images and omics.
@@ 1 Introduction
We make two contributions to this literature. First, we develop broad semi-parametric theory for network regression, showing how to incorporate node embeddings into linear regression. These network regression models are useful for causal inference on networks, as they allow practitioners to account for homophily, but they are also of substantial independent interest as an observation model. Similar models have previously appeared in Li et al. (2019); Le and Li (2022); Fosdick and Hoff (2015); He and Hoff (2019) and in concurrent work Nath et al. (2023); Chang and Paul (2024). Compared to these approaches, we require substantially weaker assumptions on the edge distribution in the network and the error distribution in the regression models. We allow weighted (i.e., real-valued) edges and allow edges to be observed with noise. Edges in the network and regression errors can be sampled from any distribution that satisfies a sub-gamma tail bound. Further, our regression models allow for heteroscedastic errors. Altogether, our results show estimating network embeddings with principal components analysis, together with ordinary least squares to estimate regression coefficients, is applicable under broad semi-parametric conditions. This contrasts with previous approaches that have largely focused on specific parametric models.
Our second contribution is to characterize how latent node-level variables (i.e., embeddings) can act as causal mediators (Imai et al., 2010). To date, causal methods for network data have mainly focused on homophily as a confounding factor. In contrast, we articulate the causal mechanisms involved when treatments influence latent social group membership. This is perhaps best illustrated by an example. Suppose we are interested in understanding and reducing adolescent smoking (Di Maria et al., 2022; Michell and Amos, 1997). Understanding why adolescents smoke is challenging, as smoking is both a sexually differentiated behavior and a social behavior. To study the effect of sex on adolescent smoking, we decompose the effect of sex into a direct effect independent of the network and an indirect effect that operates through the network. For instance, sex may directly cause higher cigarette consumption through sexed expectations about smoking. Sex may also have an indirect effect mediated by the network: adolescents may prefer to form sex-homophilous friendships, and social norms about the acceptability of smoking may vary across friend groups. That is, sex may influence an adolescent’s social circumstances, which in turns causes a change in smoking behavior. These effects correspond to natural direct and indirect effects, specialized to the network setting by treating latent community memberships as mediators.
@@
While the product-of-coefficient estimator is familiar to many, it relies on strong functional form assumptions. We want to emphasize that our characterization of latent mediation is valuable even to readers who suspect these assumptions are too strong to hold in practice. Following McFowland and Shalizi (2021), practitioners and theorists are increasingly using linear regressions for causal inference on networks, and careful charactizations of the causal structure of these models is increasingly important (Paul et al., 2022a; McFowland and Shalizi, 2021; Veitch et al., 2019, 2020; Louizos et al., 2017; Guo et al., 2020; Chu et al., 2021; Guo et al., 2020; Cristali and Veitch, 2022; Chen et al., 2022; Ogburn et al., 2022; Liu and Tchetgen Tchetgen, 2022; O’Malley et al., 2014; Leung, 2019; Egami and Tchetgen Tchetgen, 2021; Ogburn et al., 2022, 2020). In our development of the product-of-coefficients estimator, we correct the misconception that homophily can only have a confounding effect, we describe the over-control bias that can be induced by causal misspecification, and we document a network dataset where causal misspecification leads to substantially misleading results. Indeed, this work was originally motivated by a misinterpretation of coefficients in a principal components network regression model.
We expect our estimators to have significant causal and non-causal applications to outside of the network setting. Similar methods for low-rank data have been employed, for example, on spatial networks (Gilbert et al., 2021; Tiefelsdorf and Griffith, 2007; Doreian, 1981; Ord, 1975), text data (Keith et al., 2021; Veitch et al., 2020; Gerlach et al., 2018), psychometric surveys (Freier et al., 2022; Thurstone, 1947), imaging data (Zhao et al., 2020; Levin et al., 2022), and omics panels (Listgarten et al., 2010; Alter et al., 2000). Indeed, we apply our method to psychometric data, in addition to our running social network example. In data applications, we find that adolescent girls end up smoking more than adolescent boys primarily due to an indirect network effect. This suggests that public health interventions might want to disrupt or alter social group formation. In an application to psychometrics data, we find that a meditation app reduces anxiety primarily by helping study participants defuse (i.e., step back and examine) from their emotions and by helping them feel less lonely. Our findings suggest that the meditation program might improve mental health outcomes by replacing meditation modules aimed at increasing a sense of meaning in life with additional modules targeting defusion or loneliness.
@@ 2 The causal structure of mediation in latent spaces
A central goal of this paper is to formalize and estimate natural direct and indirect effects as mediated by latent positions in a network. Briefly, the idea is that node-level treatments can effect formation of social groups, and that membership in social groups can further influence nodal outcomes.
@@ 2.1 Motivating example: social and non-social elements of teenage smoking
Let us return to the adolescent smoking example mentioned previously, which is based on the Teenage Friends and Lifestyle Study, reported in Michell and West (1996), Michell and Amos (1997), Michell (1997), and Michell (2000b). The Teenage Friends and Lifestyle Study collected three waves of survey data in a secondary school in Glasgow, beginning in January 1995. Students in the study filled out a questionnaire about their lifestyle and risk-taking behaviors, including alcohol, tobacco and drug use, and additionally were asked to list six of their friends. Beyond this quantitative data, researchers also conducted in-depth qualitative work, investigating the social dynamics of the school through focus groups, classroom observations, and interviews.
In the Glasgow data, as in prior investigations, researchers found that adolescent girls smoked more than adolescent boys. Michell and West (1996) and Michell (2000b) proposed two distinct effects on smoking: one about athletic expectations, and the other about social influence. Michell (2000b) proposed that athletic expectations affected smoking as follows: both boys and girls were subject to general societal pressure to smoke, but adolescent boys were subject to additional pressure to be good at sports (Michell and West, 1996). Since smoking was generally accepted as reducing athletic performance, this influenced adolescent boys to abstain from smoking in favor of pursuing athletic social capital (Michell, 2000b).
The second proposed effect was a social effect. Smoking is known to be a collective behavior, where group members typically all partake or all abstain (Michell, 2000b). Researchers found evidence of this form of group decision-making for various risk-taking behaviors in the Glasgow study, including tobacco consumption. This group-level decision-making led to differentiated behavior amongst adolescent boys and girls because adolescent friendships are highly sex-homophilous until the onset of puberty (Mehta and Strough, 2009). Taken all together, this meant that the effect of sex on smoking was mediated by friend group membership: adolescent friend group membership was heavily determined by sex, and smoking was heavily determined by friend group. More precisely, Michell (2000b) found that smoking was mostly concentrated in friend groups composed of popular girls, unpopular students, and trouble-makers: “risk taking behaviour was heavily polarized within social categories so that, for instance, groups of individuals (and their peripherals) were in general either risk-taking or non-risk-taking […] generally groups (and their peripherals) were either all boys or all girls.”
@@ 2.3 Semi-parametric latent space structure
Lastly, under the sub-gamma model, the latent positions $X$ are sufficient for $A$. That is, the nodal covariates $T_{i},C_{i\cdot}$ and $Y_{i}$ do not directly influence the formation of edges of $A_{ij}$. The covariates $T_{i}$ and $C_{i\cdot}$ can influence edge formation, but only via the intermediary $X$. Relaxing this constraint is an interesting topic for future work, but requires different estimators for the latent positions $X$ than those we consider here (Mele et al., 2022; Binkiewicz et al., 2017).
@@ 2.4 Causal mechanisms in latent space
To develop intuition for interventions on the latent positions, consider a degree-corrected mixed membership stochastic blockmodel with five blocks, where nodes are primarily members of a single block and there is a small amount of degree heterogeneity. Interventions can increase or decrease partipication in particular communities. For instance, in Figure 3, we visualize an intervention that decreases participation in the second community while increasing participation in the third community, in the latent space.
@@
This synthetic example is primarily to develop intuition about the latent positions, and downstream consequences of an intervention. Even with this intuition, it may be difficult to understand what it means to intervene on a latent position. In practice, we believe that the best way to resolve this ambiguity is by interpreting the latent positions in the context of the relevant data application. In practice, we compute estimates of the latent positions using the adjacency spectral embedding (see Section 3). With estimates of the latent positions in hand, it is often much easier to reason about causal mechanisms, because the estimated positions can be given a concrete interpretation via inspection, auxiliary data and domain knowledge. In the psychometric data application of Section 5.2, for instance, the latent positions have a very clear interpretation. To interpret the latent positions, we find that Gaussian mixture models and varimax rotation are often valuable tools (Rohe and Zeng, 2023; Rubin-Delanchy et al., 2022; Zhang et al., 2021; Priebe et al., 2019).
To summarize: we believe that the degree-corrected mixed-memberhip stochastic blockmodel offers compelling intuition about the structures that latent positions can represent. This intuition is helpful to determine when a latent mediation model may be applicable: for instance, in settings where one suspects that social groups act as mediators. However, there is an additional confirmatory step necessary for empirical work. It is critical to confirm that the estimated latent positions do in fact capture the hypothesized mediating constructs. Estimates of latent positions might reflect other network structure unrelated to hypothesized mediators. This is a weakness of our smoking example: we know from extensive sociology research that the adolescent social network is composed of numerous small social groups. However, we cannot confirm that the social structure observed by sociologists is the same as the social structure captured by the network principal components, an issue that we discuss in more detail in Section 5.1. In contrast, in the psychometric example of Section 5.2, the estimated latent positions are clearly measures of latent constructs of interest.
@@ 2.5 Relation to peer effects
$\mathbb{E}\left[Y_{i}\,\middle|\,T_{i},C_{i\cdot},X,A,Y_{-i}\right]=\beta_{0}+T_{i}\beta_{\text{t}}+C_{i\cdot}\beta_{\text{c}}+X_{i\cdot}\beta_{\text{x}}+\beta_{\text{Ay}}\sum_{j\neq i}^{n}\frac{A_{ij}}{d_{i}}Y_{j}+\beta_{\text{At}}\sum_{j\neq i}^{n}\frac{A_{ij}}{d_{i}}T_{j}$
where $d_{i}=\sum_{j}A_{ij}$ is the degree of node $i$ (Bramoullé et al., 2009). In the linear-in-means approach, $\beta_{\text{x}}$, $\beta_{\text{Ay}}$ and $\beta_{At}$ are identified, even in the asymptotic limit, under some non-trivial assumptions about the network structure. Recent work has considered these models from a statistical point of view (Chang and Paul, 2024; Lee, 2002; Paul et al., 2022b; Trane, 2023), although we believe there are still substantial identifications concerns to address, from both a causal and a statistical perspective.
@@ 3.1 Principal components network regression
The spectral embeddings $\widehat{X}$ converge to the true $X$ uniformly over the rows of $X$, up to orthogonal non-identifiability (Levin et al., 2022; Lyzinski et al., 2014), suggesting that ordinary least squares estimates based on $\widehat{X}$ rather than $X$ may have nice asymptotic behavior. We thus construct least squares estimators for the regression coefficients as follows.

### v4

@@ Abstract
We develop a method to decompose causal effects on a social network into an indirect effect mediated by the network, and a direct effect independent of the social network. To handle the complexity of network structures, we assume that latent social groups act as causal mediators. We develop principal components network regression models to differentiate the social effect from the non-social effect. Fitting the regression models is as simple as principal components analysis followed by ordinary least squares estimation. We prove asymptotic theory for regression coefficients from this procedure and show that it is widely applicable, allowing for a variety of distributions on the regression errors and network edges. We carefully characterize the counterfactual assumptions necessary to use the regression models for causal inference, and show that current approaches to causal network regression may result in over-control bias. The method is very general, so that it is applicable to many types of structured data beyond social networks, such as text, areal data, psychometrics, images and omics.
@@ 1 Introduction
We make two contributions to this literature. First, we develop broad semi-parametric theory for network regression, showing how to incorporate node embeddings into linear regression. These network regression models are useful for causal inference on networks, as they allow practitioners to account for homophily, but they are also of substantial independent interest as an observation model. Similar models have previously appeared in Li et al. (2019); Le and Li (2022); Fosdick and Hoff (2015); He and Hoff (2019) and in concurrent work (Nath et al., 2023; Chang and Paul, 2024). Compared to these approaches, we require substantially weaker assumptions on the edge distribution in the network and the error distribution in the regression models. We allow weighted (i.e., real-valued) edges and allow edges to be observed with noise. Edges in the network and regression errors can be sampled from any distribution that satisfies a sub-gamma tail bound. Further, our regression models allow for heteroscedastic errors. Altogether, our results show that estimating network embeddings with principal components analysis, together with ordinary least squares to estimate regression coefficients, is applicable under broad semi-parametric conditions. This contrasts with previous approaches that have largely focused on specific parametric models.
Our second contribution is to characterize how latent node-level variables (i.e., embeddings) can act as causal mediators (Imai et al., 2010). To date, causal methods for network data have mainly focused on homophily as a confounding factor. In contrast, we articulate the causal mechanisms involved when treatments influence latent social group membership. This is perhaps best illustrated by an example. Suppose we are interested in understanding and reducing adolescent smoking (Di Maria et al., 2022; Michell and Amos, 1997). Understanding why adolescents smoke is challenging, as smoking is both a sexually differentiated behavior and a social behavior. To study the effect of sex on adolescent smoking, we decompose the effect of sex into a direct effect independent of the social network among adolescents and an indirect effect that operates through the network. For instance, sex may directly cause higher cigarette consumption through sex-specific expectations about smoking. Sex may also have an indirect effect mediated by the social network: adolescents may prefer to form sex-homophilous friendships, and social norms about the acceptability of smoking may vary across friend groups. That is, sex may influence an adolescent’s social circumstances, which in turns causes a change in smoking behavior. These effects correspond to natural direct and indirect effects, specialized to the network setting by treating latent community memberships as mediators.
@@
While the product-of-coefficient estimator is familiar to many, it relies on strong functional form assumptions. We want to emphasize that our characterization of latent mediation is valuable even to readers who suspect these assumptions are too strong to hold in practice. Following McFowland and Shalizi (2021), practitioners and theorists are increasingly using linear regressions for causal inference on networks, and careful characterizations of the causal structure of these models is increasingly important (Paul et al., 2022a; McFowland and Shalizi, 2021; Veitch et al., 2019, 2020; Louizos et al., 2017; Guo et al., 2020; Chu et al., 2021; Guo et al., 2020; Cristali and Veitch, 2022; Chen et al., 2022; Ogburn et al., 2022; Liu and Tchetgen Tchetgen, 2022; O’Malley et al., 2014; Leung, 2019; Egami and Tchetgen Tchetgen, 2021; Ogburn et al., 2022, 2020). In our development of the product-of-coefficients estimator, we correct the misconception that homophily can only have a confounding effect, we describe the over-control bias that can be induced by causal misspecification, and we document a network data set where causal misspecification leads to substantially misleading results. Indeed, this work was originally motivated by a misinterpretation of coefficients in a principal components network regression model.
We expect our estimators to have significant causal and non-causal applications beyond network setting. Similar methods for low-rank data have been employed, for example, on spatial networks (Gilbert et al., 2021; Tiefelsdorf and Griffith, 2007; Doreian, 1981; Ord, 1975), text data (Keith et al., 2021; Veitch et al., 2020; Gerlach et al., 2018), psychometric surveys (Freier et al., 2022; Thurstone, 1947), imaging data (Zhao et al., 2020; Levin et al., 2022), and omics panels (Listgarten et al., 2010; Alter et al., 2000). Indeed, in the sequel, we apply our method to psychometric data, in addition to our running adolescent smoking example. In data applications, we find that adolescent girls end up smoking more than adolescent boys primarily due to an indirect network effect. This suggests that public health interventions might want to disrupt or alter social group formation. In an application to psychometrics data, we find that a meditation app reduces anxiety primarily by helping study participants defuse (i.e., step back and examine) from their emotions and by helping them feel less lonely. Our findings suggest that the meditation program might improve mental health outcomes by replacing meditation modules aimed at increasing a sense of meaning in life with additional modules targeting defusion or loneliness.
@@ 2 The Causal Structure of Mediation in Latent Spaces
A central goal of this paper is to formalize and estimate natural direct and indirect effects as mediated by latent positions in a network. Briefly, the idea is that node-level treatments can affect formation of social groups, and that membership in social groups can further influence nodal outcomes.
@@ 2.1 Motivating Example: Social and Non-Social Elements of Teenage Smoking
Let us return to the adolescent smoking example mentioned previously, which is based on the Teenage Friends and Lifestyle Study, reported in Michell and West (1996), Michell and Amos (1997), Michell (1997), and Michell (2000). The Teenage Friends and Lifestyle Study collected three waves of survey data in a secondary school in Glasgow, beginning in January 1995. Students in the study filled out a questionnaire about their lifestyle and risk-taking behaviors, including alcohol, tobacco and drug use, and additionally were asked to list six of their friends. Beyond this quantitative data, researchers also conducted in-depth qualitative work, investigating the social dynamics of the school through focus groups, classroom observations, and interviews.
In the Glasgow data, as in prior investigations, researchers found that adolescent girls smoked more than adolescent boys. Michell and West (1996) and Michell (2000) proposed two distinct effects on smoking: one about athletic expectations, and the other about social influence. Michell (2000) proposed that athletic expectations affected smoking as follows: both boys and girls were subject to general societal pressure to smoke, but adolescent boys were subject to additional pressure to be good at sports (Michell and West, 1996). Since smoking was generally accepted as reducing athletic performance, this influenced adolescent boys to abstain from smoking in favor of pursuing athletic social capital (Michell, 2000).
The second proposed effect was a social effect. Smoking is known to be a collective behavior, where group members typically all partake or all abstain (Michell, 2000). Researchers found evidence of this form of group decision-making for various risk-taking behaviors in the Glasgow study, including tobacco consumption. This group-level decision-making led to differentiated behavior among adolescent boys and girls because adolescent friendships are highly sex-homophilous until the onset of puberty (Mehta and Strough, 2009). Taken all together, this meant that the effect of sex on smoking was mediated by friend group membership: adolescent friend group membership was heavily determined by sex, and smoking was heavily determined by friend group. More precisely, Michell (2000) found that smoking was mostly concentrated in friend groups composed of popular girls, unpopular students, and trouble-makers: “risk taking behaviour was heavily polarized within social categories so that, for instance, groups of individuals (and their peripherals) were in general either risk-taking or non-risk-taking […] generally groups (and their peripherals) were either all boys or all girls.”
@@ 2.3 Semi-Parametric Latent Space Structure
Lastly, under the sub-gamma model, the latent positions $X$ are sufficient for $A$. That is, the nodal covariates $T_{i},C_{i\cdot}$ and $Y_{i}$ do not directly influence the formation of edges of $A_{ij}$. The covariates $T_{i}$ and $C_{i\cdot}$ can influence edge formation, but only via the intermediary $X$. Relaxing this constraint is an interesting topic for future work, but requires different estimators for the latent positions $X$ than those we consider here (Mele et al., 2023; Binkiewicz et al., 2017).
@@ 2.4 Causal Mechanisms in Latent Space
To develop intuition for interventions on the latent positions, consider a degree-corrected mixed membership stochastic blockmodel with five blocks, where nodes are primarily members of a single block and there is a small amount of degree heterogeneity. Interventions can increase or decrease participation in particular communities. For instance, in Figure 3, we visualize an intervention that decreases participation in the second community while increasing participation in the third community, in the latent space.
@@
This synthetic example is primarily to develop intuition about the latent positions, and downstream consequences of an intervention. Even with this intuition, it may be difficult to understand what it means to intervene on a latent position. In practice, we believe that the best way to resolve this ambiguity is by interpreting the latent positions in the context of the relevant data application. In practice, we compute estimates of the latent positions using the adjacency spectral embedding (see Section 3; one can think of $A$ as a collection of proxy measurements that we use to infer $X$). With estimates of the latent positions in hand, it is often much easier to reason about causal mechanisms, because the estimated positions can be given a concrete interpretation via inspection, auxiliary data and domain knowledge. In the psychometric data application of Section 5.2, for instance, the latent positions have a very clear interpretation. To interpret the latent positions, we find that Gaussian mixture models and varimax rotation are often valuable tools (Rohe and Zeng, 2023; Rubin-Delanchy et al., 2022; Zhang et al., 2021; Priebe et al., 2019).
To summarize: we believe that the degree-corrected mixed-membership stochastic blockmodel offers compelling intuition about the structures that latent positions can represent. This intuition is helpful to determine when a latent mediation model may be applicable: for instance, in settings where one suspects that social groups act as mediators. However, there is an additional confirmatory step necessary for empirical work. It is critical to confirm that the estimated latent positions do in fact capture the hypothesized mediating constructs. Estimates of latent positions might reflect other network structure unrelated to hypothesized mediators. This is a weakness of our smoking example: we know from extensive sociology research that the adolescent social network is composed of numerous small social groups. However, we cannot confirm that the social structure observed by sociologists is the same as the social structure captured by the network principal components, an issue that we discuss in more detail in Section 5.1. In contrast, in the psychometric example of Section 5.2, the estimated latent positions are clearly measures of latent constructs of interest.
@@ 2.5 Relation to Peer Effects
$\mathbb{E}\left[Y_{i}\,\middle|\,T_{i},C_{i\cdot},X,A,Y_{-i}\right]=\beta_{0}+T_{i}\beta_{\text{t}}+C_{i\cdot}\beta_{\text{c}}+X_{i\cdot}\beta_{\text{x}}+\beta_{\text{Ay}}\sum_{j\in\mathcal{N}(i)}\frac{A_{ij}}{d_{i}}Y_{j}+\beta_{\text{At}}\sum_{j\in\mathcal{N}(i)}\frac{A_{ij}}{d_{i}}T_{j}$
where $\mathcal{N}(i)=\{j\in[n]:A_{ij}=1\}$ denotes the neighbors of node $i$ and $d_{i}=\sum_{j}A_{ij}$ is the degree of node $i$ (Bramoullé et al., 2009). In the linear-in-means approach, $\beta_{\text{x}}$, $\beta_{\text{Ay}}$ and $\beta_{At}$ are identified, even in the asymptotic limit, under some non-trivial assumptions about the network structure. Recent work has considered these models from a statistical point of view (Chang and Paul, 2024; Lee, 2002; Paul et al., 2022b; Trane, 2023), although we believe there are still substantial identifications concerns to address, from both a causal and a statistical perspective.
@@ 3.1 Principal Components Network Regression
The spectral embeddings $\widehat{X}$ converge to the true $X$ uniformly over the rows of $X$, up to orthogonal non-identifiability (Levin et al., 2022; Lyzinski et al., 2014). We thus construct least squares estimators for the regression coefficients based on $\widehat{X}$ rather than $X$, as follows.